OpenStudy (df001):

Sin(x)/1+sin(x) - sin(x)/1-sin(x) = -2tan^2(x)

OpenStudy (df001):

\[\frac{ Sin x }{ 1+\sin x } -\frac{ \sin x }{ 1 - \sin x }\]